OpenStudy (anonymous):

how do you find the inverse of f(x)=3x^2-5

OpenStudy (shamim):

Let y=f(x) Or, x=f^-1(y) Right?

OpenStudy (shamim):

Now we can write y=3x^2-5 x=?

OpenStudy (anonymous):

Simply put, you just substitute y for x, basically flipping the equation; as an example: y = 2x would invert to become x = 2y which is y = x/2. In your situation you have the function notation f(y) = something with x, which can be written as y = some thing with x the simplify this process as the guy above or possible below me said. That said, inverse function have a certain attribute which is, (f^-1(x) o f(x)) = (f(x) o f^-1(x)) = x, so you can check your work by looking making sure that evaluates properly.